OpenStudy (mathmath333):

In a mist, a runner was walking at 2 km/h, a bike passes him in the same direction. The runner could see the bike for 6 minutes and it was visible to him upto the distance of 0.6 km. the speed of the bike in km/hr was ?

OpenStudy (anonymous):

Lets call relative speed (km/h) of the bike to runner \(v\). And it's real speed \(V\). The relation between this two is: \(v=V-2\). Sincé the bike moved distance 0.6 km in 6/60 = 1/6 hours. It's speed relative to runner, \(v\), was 0.6*6=3.6 km/h. To find it's real speed use the relation above.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} v& =V-2\\ V& =v+2=3.6+2=5.6\quad km/hr\end{align}}\) unfortunately this is wrong

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} a.) 30\quad \text{km/hr}\hspace{1.5em}\\~\\ b.) 15\quad \text{km/hr}\hspace{1.5em}\\~\\ c.) 20\quad \text{km/hr}\hspace{1.5em}\\~\\ d.) 8\quad \text{km/hr}\hspace{1.5em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

rel vel V-2 so 0.6=(V-2)*6/60 so d

OpenStudy (mathmath333):

haha yep! correct !